# Feature Importance-Aware Deep Joint Source-Channel Coding for Computationally Efficient and Adjustable Image Transmission

### v1

@@ III-C Training and Loss Function Design
Let $\{\mathbf{x}^{i}\}_{i=1}^{N}$ denote an image set of batch size $N$, where $\mathbf{x}^{i}$ represents the $i$-th image. To obtain the desired importance score, we train our FAJSCC using the following feature importance-aware loss function, denoted as $\mathcal{L}_{{\textrm{FA}}}$:
@@
$\mathcal{L}_{{\textrm{FA}}}$ $:=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{{\textrm{distortion}}}(\mathbf{x}^{i},\hat{\mathbf{x}}^{i})+\eta\left\lVert\gamma_{{\textrm{train}}}-0.5(\gamma^{g}_{e}+\gamma^{g}_{d})\right\rVert_{2}^{2},$ (1)
$\gamma^{g}_{e}$ $=\frac{1}{N}\sum_{i=1}^{N}\gamma^{g,i}_{e},~{}~{}\gamma^{g}_{d}=\frac{1}{N}\sum_{i=1}^{N}\gamma^{g,i}_{d},~{}~{}\gamma_{{\textrm{train}}}<1,$
where $\eta$ is the sampling weight, $\gamma_{{\textrm{train}}}$ is the target average importance ratio in the training phase, and $\gamma^{g,i}_{e}$ and $\gamma^{g,i}_{d}$ represent the mean ratios of important features sampled for the $i$-th image by Gumbel-Softmax in the encoder and decoder, respectively. The first term $\mathcal{L}_{{\textrm{distortion}}}(\mathbf{x},\hat{\mathbf{x}})$ is an image distortion loss between original image $\mathbf{x}$ and reconstructed image $\hat{\mathbf{x}}$, as will be detailed in Section IV-A. To minimize both $\mathcal{L}_{{\textrm{distortion}}}(\mathbf{x},\hat{\mathbf{x}})$ and $\left\lVert\gamma_{{\textrm{train}}}-0.5(\gamma^{g}_{e}+\gamma^{g}_{d})\right\rVert_{2}^{2}$ in $\mathcal{L}_{{\textrm{FA}}}$, the importance scores should be high for features that contribute significantly to performance improvement and low for those with limited impact when processed via intensive computation.
@@
As a result, minimizing $\mathcal{L}_{{\textrm{FA}}}$ with Gumbel-Softmax sampling enables FAJSCC to efficiently allocate computational resources, effectively reducing the distortion between $\mathbf{x}$ and $\hat{\mathbf{x}}$. As the Gumbel-Softmax method samples important features with probabilities proportional to their importance scores, $\gamma^{g,i}_{e}$ and $\gamma^{g,i}_{d}$ fluctuate randomly for each $i$-th image transmission while keeping the batch-wise averages $\gamma^{g}_{e}$ and $\gamma^{g}_{d}$ stable during training. This enables FAJSCC to adapt to various computational budgets while ensuring stable updates during training. Another benefit of using Gumbel-Softmax is that it allows independent control of the importance ratios for the encoder and decoder during the test phase. This plays a crucial role in the experiment that measures the impact of computational complexity, which will be discussed later.
@@ IV-A Experimental Setting
Baselines and Architectures: We implement two versions of FAJSCC: one as illustrated in Figure 2 with details in Section III-B (FAJSCC w/ SA), and another without SNR adaptation (FAJSCC w/o SA) where the two-step SNR feature extraction layers are deleted. The other deepJSCC baselines are ranging from the original deepJSCC [6] to the recent SwinJSCC [8]. To differentiate between these baselines, we refer to the initial deepJSCC [6] based on the convolution blocks and the deepJSCC based on residual convolution blocks [7] as ConvJSCC and ResJSCC, respectively.
We implement our FAJSCCs (FAJSCC w/ SA and FAJSCC w/o SA) and other baselines with different network scales to evaluate their performance under varying computational resources. For ConvJSCC and ResJSCC, the feature channel sizes are set to $32$ and $64$ for small and base models, respectively. For SwinJSCC, the feature channel sizes at each stage $[C_{1},C_{2},C_{3},C_{4}]$ are set to $[40,60,80,160]$ for the small model and $[60,90,120,200]$ for the base model. For FAJSCCs, the feature channel sizes at each stage $[C_{1},C_{2},C_{3},C_{4}]$ are set to $[40,60,80,260]$ for the small model and $[60,90,120,360]$ for the base model. For FAJSCCs and SwinJSCC, the number of main computational blocks per stage (i.e., FA block and Swin transformer block, respectively) is set to $2$.
Datasets and Training To train models, we use the DIV2K training dataset [45], which consists of $800$ high-resolution images. To maintain consistent image resolution within each batch, we randomly crop images to a resolution of $128\times 128$ with a batch size of $32$ during training. The image distortion loss $\mathcal{L}_{{\textrm{distortion}}}$ is set to MSE and $1-{\textrm{SSIM}}$ when evaluating models for PSNR and SSIM, respectively. The baseline models are trained to minimize $\mathcal{L}_{{\textrm{distortion}}}$ while our FAJSCCs are trained to minimize $\mathcal{L}_{{\textrm{FA}}}$ in (1). The hyperparameter settings of $\mathcal{L}_{{\textrm{FA}}}$ are $\eta=1.0$ and $\gamma_{{\textrm{train}}}=0.5$. Note that $\gamma_{{\textrm{train}}}$ is the target average importance ratio in training phase; the importance ratio during test phase could be different, e.g., Figure 11. All models are trained using Adam optimizer with a learning rate $0.0001$ for $200$ epochs.
We consider three CPP values $[\frac{1}{12},\frac{1}{16},\frac{1}{24}]$ and four SNR values $[1,4,7,10]$ dB under AWGN and fast Rayleigh fading channels. For all models except FAJSCC w/ SA, the training SNR is the same as the test SNR. In contrast, FAJSCC w/ SA is trained with SNR values uniformly sampled from $[1,4,7,10]$ dB. Since we observed unstable updates in ConvJSCC and ResJSCC during training with small CPP values, we selected the best-performing model out of five trained models for each to ensure reliable results. Test results are obtained from the DIV2K validation dataset [45] for numerical evaluation and the Kodak dataset [46] for visual inspection in Figure 12.
@@
Figure 9: Efficiency comparison in GFLOPs and memory for small and base models under AWGN channel at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=7$dB.
@@ IV-B Main Result
Architecture Efficiency: Figure 9 shows PSNR performance in terms of computational cost (GFLOPs, solid curves) and memory usage (MB, dotted curves) across different model scales. Here, GFLOPs are measured when transmitting $512\times 768$ resolution image. In this figure, our FAJSCCs demonstrate high efficiency: Regarding computational cost, FAJSCCs achieve the highest PSNR while using smaller computational resources than SwinJSCC across different scales. Regarding memory usage, FAJSCCs achieve the highest PSNR while using smaller memory than SwinJSCC across different scales. This performance gain primarily comes from FAJSCC’s importance-wise feature processing, as well as two key techniques: (1) leveraging feature correlations through depth-wise convolution to mitigate channel errors and (2) using GDF to filter out unimportant or noisy information. These techniques collectively contribute to two fundamental aspects of communication—error correction in channel coding and compression in source coding. These advantages lead FAJSCCs to outperform SwinJSCC while using smaller computational and memory resources.
PSNR and SSIM Results: Figures 8 and 10 show the PSNR and SSIM performance results across various communication environments. Our FAJSCCs outperform the previous models for different performance metrics under various communication environments. Even though our FAJSCC w/ SA is trained across multiple different SNRs, while others are trained at training SNR that exactly matches test SNR, our FAJSCC w/ SA outperforms the previous deepJSCC models. This demonstrates the efficient and SNR-adaptive abilities of FAJSCC w/ SA.
When compared to the FAJSCC w/o SA version, our FAJSCC w/ SA achieves similar or better performance at low SNRs by leveraging SNR information. This demonstrates the effectiveness of our two-step SNR feature extraction approach. Moreover, it incurs only a small extra computational cost, as shown in Figure 9, where the extra overhead due to SNR processing is nearly imperceptible. At high SNRs, FAJSCC w/ SA performs slightly worse than FAJSCC w/o SA. Since losses at low SNRs are higher than those at high SNRs, FAJSCC w/ SA trained on random SNRs prioritizes minimizing losses at low SNRs. This trend aligns with findings from other SNR-adaptive deepJSCC models [8, 13].
Computation Resource Adjustment: Our FAJSCC is the first deepJSCC model capable of dynamically adjusting computational resources by modifying the importance ratio. This adjustment can be applied independently to the encoder and decoder by varying the importance ratio $\gamma_{e}$ for the encoder’s FA blocks and $\gamma_{d}$ for the decoder’s FA blocks, without needing to agree with each other’s value. Since available computational resources for the encoder and decoder may vary over time, this flexible adjustment mechanism enhances the practicality of FAJSCC for real-world communication applications.
Figure 11 illustrates PSNR performance versus GFLOPs as the importance ratios vary under different communication environments. The values of $\gamma_{e}$ and $\gamma_{d}$ range from $[0.0,0.2,0.4,0.5,0.6,0.8,1.0]$, and is fixed at $0.5$ when unspecified. As shown, FAJSCC outperforms SwinJSCC, demonstrating high performance across various computational resources and communication environments. Except for the case of varying $\gamma_{e}$ where curves are nearly flat, PSNR improves significantly as the importance ratio increases from 00 to $0.5$. However, beyond $0.5$, the performance gain diminishes or even the performance slightly decreases as the importance ratio approaches to $1.0$. This indicates that less important features do not require extensive computational resources, validating our approach of processing important and less important features differently for improved efficiency.
Complexity Demands at Encoder and Decoder: Recalling that deepJSCC involves the joint optimization of source coding and channel coding, we can further investigate the image reconstruction performance from three different perspectives. The first perspective is the encoder for source and channel coding where the extracted features should be both concise and robust to channel noise. The second perspective is the decoder for source coding where images are regenerated. The final perspective is the decoder for channel coding where error correction is performed at the deepJSCC decoder. By considering these components, our experiments with varying values of $\gamma_{e}$ and $\gamma_{d}$ reveal the relationship between computational complexity and the contribution of each component to reconstruction performance. It is important to note that all previous deepJSCC models have fixed computational complexity at encoder and decoder and cannot adjust computational resources without modifying network scales or architectures [6, 8, 13]. As a result, such a discussion has not been possible in previous models.
First, when varying $\gamma_{e}$ (triangles in Figure 11), no significant change in PSNR is observed at both ${\textrm{SNR}}=1$dB (low) and ${\textrm{SNR}}=10$dB (high). This indicates that robust and concise feature extraction in the encoder does not require substantial computational resources, and adjusting the encoder’s computational budget has minimal impact on overall communication performance. On the other hand, when varying $\gamma_{d}$ (crosses in Figure 11), PSNR exhibits different behaviors depending on the SNR. The PSNR variation at ${\textrm{SNR}}=1$dB is approximately $0.4$dB, which is twice of the $0.2$dB variation at ${\textrm{SNR}}=10$dB. Since the image regeneration tasks are the same at both SNRs and thus the required computations are also the same, we can conclude that error correction at the decoder demands significant computational complexity. Our findings suggest that the decoder plays a more critical role than the encoder in deepJSCC performance, especially when SNR is low. This highlights the need to reconsider the conventional symmetric deepJSCC design, where the encoder and decoder typically have symmetric structures and computational resources [6, 8, 13].
@@
Figure 12: The first row shows the transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=7$dB under the AWGN channel, and the third row shows the transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=1$dB under the AWGN channel. The second and fourth rows show the detailed result of the green box in the original images of the first and third rows.
@@
Visual Inspection: Figure 12 visualizes two selected images from Kodak dataset. The first row shows a transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=7$dB under the AWGN channel, and the third row shows a transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=1$dB under the AWGN channel. The second and fourth rows provide zoomed-in views of the green box in the original images of the first and third rows.
@@
As illustrated in the figure, FAJSCCs reconstruct image details with fewer artifacts and greater clarity compared to previous deepJSCC methods. For example, in the second row, our FAJSCCs can reconstruct the ship’s sail more clearly than others. In the fourth row, our FAJSCCs can reconstruct the water hole more clearly than others. These qualitative results further validate the superior image transmission capabilities of FAJSCCs, aligning with the quantitative performance results. Additionally, as previously discussed, FAJSCC w/o SA often outperforms FAJSCC w/ SA at high SNR (e.g., the first two rows at $7$dB). This is because FAJSCC w/ SA, trained at randomly sampled SNR, tends to prioritize performance at low SNRs. At low SNR (e.g., last two rows at $1$dB), FAJSCC w/ SA outperforms its w/o SA version.

### v2

@@ III-D Training and Loss Function Design
To encourage the window importance head to assign a high score to $z_{{\textrm{imp}}}(i,j)$ when the $(i,j)$-th window is important and a low score otherwise, we introduce a penalty term that discourages high scores for less important windows. Let $\{\mathbf{x}^{i}\}_{i=1}^{N}$ denote an image set of batch size $N$, where $\mathbf{x}^{i}$ represents the $i$-th image. To obtain the desired importance score, we train our FAJSCC using the following feature importance-aware loss function, denoted as $\mathcal{L}_{{\textrm{FA}}}$:
@@
$\mathcal{L}_{{\textrm{FA}}}$ $:=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{{\textrm{distortion}}}(\mathbf{x}^{i},\hat{\mathbf{x}}^{i})+\eta\left\lVert\gamma_{{\textrm{train}}}-\frac{(\bar{\mathbf{z}}^{e}_{{\textrm{ST}}}+\bar{\mathbf{z}}^{d}_{{\textrm{ST}}})}{2}\right\rVert_{2}^{2},$ (1)
$\bar{\mathbf{z}}^{e}_{{\textrm{ST}}}$ $=\frac{1}{N}\sum_{i=1}^{N}\bar{\mathbf{z}}^{e,i}_{{\textrm{ST}}},~~\bar{\mathbf{z}}^{d}_{{\textrm{ST}}}=\frac{1}{N}\sum_{i=1}^{N}\bar{\mathbf{z}}^{d,i}_{{\textrm{ST}}}~~\gamma_{{\textrm{train}}}<1,$
[…]
@@
As a result, minimizing $\mathcal{L}_{{\textrm{FA}}}$ with Gumbel-Softmax sampling enables FAJSCC to efficiently reduce the distortion between $\mathbf{x}$ and $\hat{\mathbf{x}}$ by appropriately sampling important windows. As the Gumbel-Softmax method samples important windows with probabilities proportional to their window importance, $\bar{\mathbf{z}}^{e,i}_{{\textrm{ST}}}$ and $\bar{\mathbf{z}}^{d,i}_{{\textrm{ST}}}$ fluctuate randomly for each $i$-th image transmission while keeping the batch-wise averages $\bar{\mathbf{z}}^{e}_{{\textrm{ST}}}$ and $\bar{\mathbf{z}}^{d}_{{\textrm{ST}}}$ stable during training. This enables FAJSCC to adapt to various computational budgets while ensuring stable updates during training. Another benefit of using Gumbel-Softmax is that it allows independent control of the importance ratios for the encoder and decoder during the test phase. This plays a crucial role in the experiment that measures the impact of computational complexity, which will be discussed later.
@@ IV-A Experimental Setting
Baselines and Architectures: We implement two versions of FAJSCC: The first is as illustrated in Figure 2 with details in Section III. The other is without the selective deformable self-attention proposed in Section III-B, i.e., unused offset and window importance heads in the attention family tree, and the selective deformable self-attention module in the main computational path are removed. We call the latter version light-attention deepJSCC (LAJSCC). LAJSCC uses only the axis-dimension specialized computation proposed in Section III-A to verify its computational efficiency. The other deepJSCC baselines are ranging from the original deepJSCC [6] to the recent SwinJSCC [8] and LICRFJSCC [32]. To differentiate between these baselines, we refer to the initial deepJSCC [6] based on the convolution blocks and the deepJSCC based on residual convolution blocks [7] as ConvJSCC and ResJSCC, respectively. The detailed implementation code is available at github.com/hansung-choi/FAJSCCv2.
Datasets and Training To train models, we use the DIV2K training dataset [51], which consists of $800$ high-resolution images. To maintain consistent image resolution within each batch, we randomly crop images to a resolution of $128\times 128$ with a batch size of $32$ during training. The image distortion loss $\mathcal{L}_{{\textrm{distortion}}}$ is set to MSE and $1-{\textrm{SSIM}}$ when evaluating models for PSNR and SSIM, respectively. Our FAJSCC is trained to minimize $\mathcal{L}_{{\textrm{FA}}}$ in (1) while others are trained to minimize $\mathcal{L}_{{\textrm{distortion}}}$. The hyperparameter settings for $\mathcal{L}_{{\textrm{FA}}}$ are $\eta=1.0$, $\gamma_{{\textrm{train}}}=0.5$ , and $\tau=1$. Note that $\gamma_{{\textrm{train}}}$ is the target average importance ratio in the training phase; the importance ratio during the test phase could be different, e.g., Figure 8. All models are trained using Adam optimizer with a learning rate $0.0001$ for $200$ epochs.
We consider four CPP values $\frac{1}{12},\frac{1}{16},\frac{1}{24},\frac{1}{32}]$ and four SNR values [1,4,7,10][1,4,7,10]dB under AWGN and fast Rayleigh fading channels. For all models, the training SNR is the same as the test SNR. Since we observed unstable updates in ConvJSCC and ResJSCC during training with small CPP values, we selected the best-performing model out of five trained models for each to ensure reliable results. Test results are obtained from the DIV2K validation dataset [51] for numerical evaluation and the Kodak dataset [52] for visual inspection in Figure 9.
@@ IV-B Main Result
TABLE I: Comparison for GFLOPs and memory usage of deepJSCCs under AWGN channel, SNR=10dB10\ $\mathrm${dB}, and CPP=$1/16$.
PSNR $\mathrm{(dB)}$ GFLOPs Memory (MB)
ConvJSCC [6] 22.50 4.48 0.64
ResJSCC [7] 25.51 5.73 0.77
SwinJSCC [8] 26.46 20.02 71.26
LICRFJSCC [32] 25.61 9.02 3.84
LAJSCC (ours) 27.36 11.35 5.27
FAJSCC (ours) 27.92 17.58 13.56
@@
Figure 9: The first row shows the transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=10$dB under the AWGN channel, and the third row shows the transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=1$dB under the AWGN channel. The second and fourth rows show the detailed results of the green box in the original images of the first and third rows.
Architecture Efficiency: Table I shows PSNR performance in terms of computational cost (GFLOPs) and memory usage (MB) under AWGN channel, SNR=1010dB, and CPP=$1/16$. Here, GFLOPs are measured when transmitting $512\times 768$ resolution image. In this table, our LAJSCC and FAJSCC achieve the highest PSNR while using smaller computational resources and memory than SwinJSCC [8].
PSNR and SSIM Results: Figures 6 and 7 show the PSNR and SSIM performance results across various communication environments. Our LAJSCC and FAJSCC outperform the previous models for different performance metrics under various communication environments. This performance gain primarily comes from our dimension-specialized computation approaches: (1) Using only axis-dimension specialized operations (LAJSCC) shows higher performance than the recent SOTA SwinJSCC in most communication environments, while requiring computational resources about half and $14$ times smaller memory. This is a notable achievement, as even the recent SOTA lightweight model LICRFJSCC [32] reduces computational resources by shrinking feature sizes when applying self-attention, which inherently limits its performance, making it impossible to reach that of SwinJSCC. (2) Upon the axis-dimension specialized operations, our proposed selective deformable self-attention (FAJSCC) even more enhances the image transmission performance by utilizing deformation and feature importance. These advantages lead FAJSCC to outperform SwinJSCC while using smaller computational and memory resources.
Computation Resource Adjustment: Our FAJSCC is the first deepJSCC model capable of independently adjusting computational resources of the encoder and decoder. It is important to note that typical previous deepJSCC models have fixed and symmetric computational complexity at encoder and decoder [6, 8, 13] or the decoder’s computational complexity is inherently constrained by feature deletion at the encoder [33, 37]. As a result, such independent computational resource adjustment has not been possible in previous models. On the contrary, FAJSCC achieves this adjustment by independently varying the importance ratio $\gamma_{e}$ for the encoder’s FA blocks and $\gamma_{d}$ for the decoder’s FA blocks, without needing to agree with each other’s value. Since available computational resources for the encoder and decoder may vary over time, this flexible adjustment mechanism enhances the practicality of FAJSCC for real-world communication applications.
Figure 8 illustrates PSNR performance versus GFLOPs as the importance ratios vary under different communication environments. The values of $\gamma_{e}$ and $\gamma_{d}$ range from $[0.0,0.2,0.4,0.5,0.6,0.8,1.0]$, and is fixed at $0.5$ when unspecified. Except for the case of varying $\gamma_{e}$ where curves are nearly flat, PSNR improves significantly as the importance ratio increases from 0 to $0.5$. However, beyond $0.5$, the performance gain diminishes or even slightly decreases as the importance ratio approaches $1.0$. This indicates that less important features do not require extensive computational resources, validating our approach of selectively enhancing important features to increase computational efficiency. Moreover, this selective feature enhancement strategy does not have any information loss for adjusting computational complexity compared to the previous resource-adjustable deepJSCC methods that delete features [33, 37]. As a result, our FAJSCC maintains higher performance for various $\gamma$ than the current SOTA SwinJSCC, which shows 25.0525.05dB PSNR at SNR=11dB and 26.8126.81dB PSNR at SNR=1010dB.
Complexity Demands at Encoder and Decoder: Recalling that deepJSCC involves the joint optimization of source coding and channel coding, we can further investigate the image reconstruction performance from three different perspectives. The first perspective is the encoder for source and channel coding where the extracted features should be both concise and robust to channel noise. The second perspective is the decoder for source coding where images are regenerated. The final perspective is the decoder for channel coding, which involves perceiving the meaning of the noisy received signal. By considering these components, our experiments with varying values of $\gamma_{e}$ and $\gamma_{d}$ reveal the relationship between computational complexity and the contribution of each component to reconstruction performance. It is important to note that previous deepJSCC models have fixed computational complexity at the encoder and decoder [6, 8, 13] or the decoder’s computational complexity is inherently constrained by feature deletion at the encoder [33, 37]. As a result, such a discussion has not been possible in previous models.
First, when varying $\gamma_{e}$ (triangles in Figure 8), no significant change in PSNR is observed at both ${\textrm{SNR}}=1$dB (low) and ${\textrm{SNR}}=10$dB (high). This indicates that robust and concise feature extraction in the encoder does not require substantial computational resources, and adjusting the encoder’s computational budget has minimal impact on overall communication performance. On the other hand, when varying $\gamma_{d}$ (crosses in Figure 8), PSNR exhibits different behaviors depending on the SNR. The PSNR variation at ${\textrm{SNR}}=1$dB is approximately 0.280.28dB, which is about three times of the 0.090.09dB variation at ${\textrm{SNR}}=10$dB. Since the image regeneration tasks are the same at both SNRs, and thus the required computations are also the same, we can conclude that perceiving the meaning of the noisy received signal at the decoder demands significant computational complexity. Our findings suggest that the decoder plays a more critical role than the encoder in deepJSCC performance, especially when SNR is low. This highlights the need to reconsider the conventional symmetric deepJSCC design, where the encoder and decoder typically have symmetric structures and computational resources [6, 8, 13]. We show that designing a larger decoder than the encoder is effective in the low SNR regime, even with a simple vanilla implementation, in the ablation studies at Section IV-C.
@@
Visual Inspection: Figure 9 visualizes two selected images from Kodak dataset. The first row shows a transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=10$dB under the AWGN channel, and the third row shows a transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=1$dB under the AWGN channel. The second and fourth rows provide zoomed-in views of the green box in the original images of the first and third rows.
@@
As illustrated in the figure, FAJSCC and LAJSCC reconstruct image details with fewer artifacts and greater clarity compared to previous deepJSCC methods. For example, in the second row, ours can reconstruct the ship’s sail more clearly than others. In the fourth row, ours can reconstruct a more detailed water hole than others. These qualitative results further validate the superior image transmission capabilities of our proposed models, aligning with the quantitative performance results.
TABLE II: The ablation study for verifying the effectiveness of our proposed methods and deepJSCC complexity analysis under AWGN channel and CPP=$1/32$. PSNR ($\mathrm{dB}$) values of each SNR, and GFLOPs and Memory (MB) usages are evaluated.
SNR 1dB 4dB 7dB 10dB GFLOPs Memory (MB)
SwinJSCC [8] 24.01 24.89 25.48 25.90 20.01 71.21
LAJSCC 23.85 24.91 25.84 26.56 11.33 5.21
FAJSCC 23.96 25.15 26.07 26.86 17.53 13.47
FAJSCC w/ PGB 23.99 25.16 26.09 26.90 17.53 13.47
FAJSCC w/o AT 23.87 25.04 26.18 27.02 18.87 14.49
FAJSCC w/o FD 24.09 25.10 25.88 26.71 16.84 13.32
FAJSCC w/o LA 23.98 24.97 25.80 26.55 17.47 13.11
LAFAJSCC 24.05 25.09 25.80 26.53 15.00 9.34
FALAJSCC 23.91 24.94 25.82 26.65 15.10 9.34
IV-C Ablation Studies
The result of the top three models in Table II shows that our proposed axis dimension specialized processing (LAJSCC) is sufficient to surpass the current SOTA SwinJSCC [8] with much smaller computational and memory resources. Moreover, using a selective deformable self-attention (FAJSCC) further improves image transmission performance for various SNR settings. In addition, we conduct extensive ablation studies, examining each component point by point to verify the effectiveness of our proposed methods.
Attention Feature Extraction: To verify the effectiveness of our proposed attention family tree structure, we compare two models, which are variants of our FAJSCC: FAJSCC with partial gradient block (FAJSCC w/ PGB) and FAJSCC without attention family tree (FAJSCC w/o AT). The architecture of FAJSCC w/ PGB is the same as the FAJSCC, but we prevent backward-gradient flow from the window importance head to the spatial importance in the attention family tree of Figure 3 during training. Since the spatial importance branch only learns to extract information for spatial attention in this training setting, the window importance head should predict window importance based only on the spatial attention information. The results show that FAJSCC w/ PGB performance is very similar to FAJSCC. This means that spatial attention and window importance are highly related, so that information for extracting spatial attention is sufficient to extract window importance. Thus, our attention family tree structure extracts each attention feature efficiently by removing redundant overlapping processing steps based on the relations of offset, attention feature, and window importance.
Another direction of the ablation study also proves the efficiency of our attention family tree. FAJSCC w/o AT is constructed by replacing the attention family tree with the independent parallel predictors for each attention feature, offset, and window importance. The results show that FAJSCC shows better or similar performance compared to the FAJSCC w/o AT with smaller computational and memory resources. This once again verifies the efficiency of our attention family tree.
Attention Mechanism: To verify the effectiveness of our feature deformation for selective self-attention, we make a new variant of FAJSCC that does not use feature deformation (FAJSCC w/o FD). The results show that FAJSCC is better than FAJSCC w/o FD in most SNR settings, which validates the need to use feature deformation before self-attention as we proposed. We also checked whether using spatial attention and channel attention mechanisms enhances feature processing ability by making a new variant of FAJSCC that does not use these lightweight attentions (FAJSCC w/o LA). The results show that FAJSCC is better than FAJSCC w/o LA in most SNR settings, which validates the need to use each lightweight attention before depthwise convolution and pointwise convolution as we proposed.
Complexity Demands of DeepJSCC: In Section IV-B, we showed that perceiving the meaning of the noisy received signal requires the highest computational resources; thus, an asymmetric structure that takes a larger decoder than the encoder will have high efficiency when SNR is low. We show this by comparing three models: 1) FAJSCC, 2) LAFAJSCC where the encoder is from LAJSCC and the decoder is from FAJSCC, and 3) FALAJSCC where the encoder is from FAJSCC and the decoder is from LAJSCC. The results show that LAFAJSCC is better or comparable to the FAJSCC in low SNR regimes while using lower computational resources. Moreover, FALAJSCC is worse than LAFAJSCC in low SNR regimes while using similar computational resources. This highlights the importance of adapting the encoder–decoder asymmetry to the channel condition. Therefore, utilizing computational complexity demands under the given SNR will lead to highly efficient communication systems.